OpenStudy (anonymous):

In solving the equation (x + 3)(x – 3) = -8, Eric stated that the solution would be x + 3 = -8 => x = -11 or (x – 3) = -8 => x = -5 However, at least one of these solutions fails to work when substituted back into the original equation. Why is that? Please help Eric to understand better; solve the problem yourself, and explain your reasoning.

OpenStudy (anonymous):

x = 1 because if the two binomials are multiplied, you get: \[x^{2}-9 = -8\] \[x^{2} = -8+9\] \[\sqrt{x}\] or x = 1 or -1